OpenStudy (anonymous):

What is the value of the expression when z = –28

OpenStudy (anonymous):

What is the value of the expression \[\frac{ z-4 }{ -8 }\] when z = –28

OpenStudy (anonymous):

@Chiko_1278

OpenStudy (anonymous):

do u have any answer choices??

OpenStudy (anonymous):

no it is a box

OpenStudy (anonymous):

well it got me -27.5...

OpenStudy (anonymous):

ok i have one moreWhat is the mean of the set of numbers? {–9, –8, 4, –1}

OpenStudy (anonymous):

but yeah is 3 i did it on paper work

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

- or +

OpenStudy (anonymous):

??

OpenStudy (anonymous):

+

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

What is the mean of the set of numbers? {–9, –8, 4, –1}

OpenStudy (pandagurl):

to find the mean, add all the numbers together, then divide them by how many of them there are

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

so what is the answer

OpenStudy (pandagurl):

so, -9 + -8 + 4 +-1= -14

OpenStudy (anonymous):

A. –5.5 B. –3.5 C. –2.8 D. 14

OpenStudy (pandagurl):

then -14 / 4= -3.5. So the Answer is B